OpenStudy (anonymous):

F(x)=2x^2-5x+6; f(x-1)

OpenStudy (freckles):

replace x with x-1

OpenStudy (michele_laino):

Hint: it is suffice to replace \(x\) with \(x-1\), so we can write: \[f\left( {x - 1} \right) = 2{\left( {x - 1} \right)^2} - 5\left( {x - 1} \right) + 6 = ...?\]